### v2

# Rethinking Reuse in Dependency Supply Chains: Initial Analysis of NPM packages at the End of the Chain

Abstract.
The success of modern software development can be largely attributed to the concept of code reuse, such as the ability to reuse existing functionality via third-party package dependencies, evident within massive package networks like NPM, PyPI and Maven. For a long time, the dominant philosophy has been to ‘reuse as much as possible, without thought for what is being depended upon’, resulting in the formation of large dependency supply chains that spread throughout entire software ecosystems. Such heavy reliance on third-party packages has eventually brought forward resilience and maintenance concerns, such as security attacks and outdated dependencies.
In this vision paper, we investigate packages that challenge the typical concepts of reuse–that is, packages with no dependencies themselves that bear the responsibility of being at the end of the dependency supply chain. By avoiding dependencies, these packages at the end of the chain may also avoid the associated maintenance and resilience risks typically beyond their control. For our initial analysis, we look at a sample of heavily-depended upon NPM packages, and identify that such end-of-chain packages make up a significant portion of these critical dependency chain (over 50%). We find that these end-of-chain packages vary in characteristics and not just packages that can be easily replaced: an active, well-maintained package at the end of the chain; a ”classical” package that has remained unchanged for 11 years; a trivial package nested deep in the dependency chain; a package that may appear trivial; and a package that bundled up and absorbed its dependencies.
We then ask ourselves: Should project maintainers follow suit and become less dependent on third-party packages? What would the impact on supply chains be if any of these critical packages were abandoned or attacked by bad actors? The vision of this paper is to advocate for a shift in software development practices toward minimizing reliance on third-party packages, particularly those at the end of dependency supply chains. We argue that these end-of-chain packages offer unique insights, as they play a key role in the ecosystem. At the same time, we do not want to erase the success of NPM reuse, but we need to understand how to strategically reuse so that the benefits are fully realized by the maintainer and the ecosystem as a whole.
1. Introduction
The practice of reusing code via third-party packages is a common part of software development, evident by the popularity of package managers such as Node Package Manager (NPM), hosting over 2.1 million packages (or packages) as of September 2022 (The OpenJS Foundation, [n. d.]). The abundance of such packages enables developers to integrate already-made solutions over writing their own implementations, with well-established benefits for productivity, code quality and defect reduction (Lim, 1994; Basili et al., 1996).
Despite the ubiquity of such code reuse, there is not a lack of risk; recent work has focused on the impact of dependencies on the security and sustainability of software projects (Decan et al., 2018; Wattanakriengkrai et al., 2022), especially in regards to resilience and maintenance. Many NPM packages depend directly on a number of other packages, that then have their own dependencies and so on, resulting in lengthy chains of indirect dependencies. Prior work has considered code reuse as a double-edged sword (Gkortzis et al., 2019): the maturity and open-source nature of these packages can make systems more secure, but at the same time the increased attack surface can introduce vulnerabilities. Maintainers and users of a package may not necessarily be aware of the full scale and impact of a package’s dependency chain - the result is opportunities for unsecured links on the supply chain to introduce security and maintenance issues to upstream dependent packages.
The unpublishing of the left-pad package in 2016 subsequently broke numerous other important packages on NPM, including software maintained by large companies such as Meta (Collins, 2016). While Meta’s React package didn’t directly depend on this trivial, 11 line package, it was impacted via its dependency chain. Increasingly, there are also social aspects such as protestware to consider - the introduction of potentially malicious code for political or financial reasons (Kula and Treude, 2022). In the case of the node-ipc package, when ran by users from Russia or Belarus, the code would attempt to replace all files with a heart emoji in protest of the war in Ukraine. In response to these concerns and high-profile cases that propagate across an entire ecosystem such as Maven’s Log4Shell vulnerability (YfryTchsGD, 2021), efforts to tighten the software supply chain have emerged. The Alpha and Omega Project specifically aims to help the most critical open source projects improve their security postures (Foundation, 2024). More and more, software developers are looking to reduce dependencies and their overhead (Lu, 2024).
Despite the popular culture of reuse that is apparent within the NPM ecosystem, some packages curiously abstain from the use of third-party dependencies themselves, making them the end of the dependency supply chain. In doing so, these packages may also be avoiding the risk factors that come with dependency use. Previous work investigating the NPM ecosystem has focused on the impact of lengthy dependency chains, but not necessarily the reverse, and the results of these studies may not generalise to the no dependency case. In this vein, there may be undiscovered benefits to refraining from code reuse in specific scenarios wherein the ‘buck’ of the supply chain stops with this package.
In this paper, we argue that analysis of software ecosystems like NPM should not ignore the role of end-of-chain packages, and how the unforeseen benefit may stack up against the productivity benefits of code reuse. For the most popular packages on NPM, our preliminary results indicate that packages with no dependencies are indeed evident in the dependency chain, comprising just over 50% of the dependency chain in our sample. We select a series of motivating cases, which we believe characterise some the different types of end-of-chain dependency packages that exist in the NPM ecosystem: popular packages, ‘classical’ packages, trivial packages, complex trivial packages and packages that absorb their dependencies. From these cases, we present a series of open research questions to guide researchers in investigating, characterising and understanding the perils and best practice of these no dependency packages, their impact and the motivations behind reducing dependencies. Our new idea is to explore how understanding the packages at the end of the chain may impact the quality and resilience of the supply chain itself, and further reveal more insights into the complicated trade-offs of code reuse in software development.
2. Extracting End of Chain from the Top 10 NPM packages
We conduct an initial analysis of the top ten most depended upon NPM packages. We begin by looking at the top 10 most depended upon packages in the NPM ecosystem and their chain of upstream dependencies. Being one of the largest and most studied software ecosystems (The OpenJS Foundation, [n. d.]; Zimmermann et al., 2019; Kula et al., 2017), we choose to limit this initial analysis to only the NPM ecosystem. We select this subset of packages because they represent a sample of the packages critical to the NPM dependency chain. First, we utilise the packages.io [^1] https://packages.io/ API to get the top 10 packages sorted by dependency count. Next, for those 10 packages, we mine the NPM registry [^2] https://registry.npmjs.org for dependency data, and recursively build their chain of dependencies; that is, the packages they depend upon. We look only at the dependencies field on the NPM registry, and ignore devDependencies and peerDependencies. devDependencies pertain to only dependencies required during development of the plug-in which are not installed in production, while peerDependencies often denotes that a package is a plug-in for another package, but it may or may not actually use APIs from that package. We limit our analysis to regular dependencies to avoid the complexity of these cases for this investigation. Additionally, no recursive dependencies are present in this dataset. Table 1 shows the final dataset statistics.
Table 1. Statistics for the total dependency chain of the top 10 most depended upon packages on NPM.
Root NPM packages 10
Dependency chain size 504
Unique packages in dependency chain 216
End-of-Chain packages 115
End-of-Chain packages (manually verified) 114
2.1. Prevalence of end-of-chain packages
Out of 216 packages within this dependency chain, we observe that 115 (53.24%) have no dependencies according to the NPM registry. Additionally, five of the top ten most depended upon packages (the roots of the dependency tree) report no dependencies via this method. In section 3 we take a deeper look at example cases from this sample; however, as part of our deeper analysis we also uncover that the NPM dependency data is not always consistent with the package.json file available on GitHub. For example, prettier reports zero dependencies, despite having dependencies listed in the package.json. However, this is the only instance in our dataset, reducing the manually verified zero dependency number to 114 packages (52.78%). The final number of end-of-chain packages in the top ten most depended upon is four: typescript, eslint-plugin-react-hooks, moment and @types/lodash. This analysis reveals that some of the most popular packages in the NPM ecosystem have no dependencies.
3. Exploring Five Cases of End-of-Chain
Based on the dataset results, we now present five motivating cases of no dependency packages at the end of the dependency chain, their impacts on the dependency chain and the implications of their lack of dependencies. We select these cases from our sample based on their prospective interest to the software engineering community. The first case describes the case of an active, well-maintained package at the end of the dependency chain. The second case describes the case of a ‘classical’ package that has remained unchanged for 11 years, that’s lack of activity may be a side effect of its simplicity and lack of dependencies. In the third case, we take a look at the case of a trivial package nested deep in the dependency chain of the popular eslint package. Forth, we observe the case of a package that may appear trivial, but package history and activity suggests otherwise. Finally, we look at the case of a package that bundled up and absorbed its dependencies.
3.1. Case 1: The popular package (Most Depended Upon)
The most depended upon NPM package according to packages.io data also has zero dependencies: the typescript package [^3] https://github.com/microsoft/TypeScript. As many projects utilize TypeScript, the popularity of this package is not surprising. The package is often installed as part of the build process for a package, where TypeScript will be transpiled into JavaScript. The package is 32MB and the GitHub repository is extremely active, with a commit in the last two days, 5000+ issues and 338 pull requests. The repository also lists 761 contributors. There is also a level of expected support, as the package is maintained by Microsoft. In terms of security, the repository has a security policy and no reported vulnerabilities on Snyk (Snyk, 2024b). A look at the history of this package’s dependencies reveals that it has always had zero dependencies.
We look at this case as an example of an active, popular package at the end of the dependency chain. Due to its position in the JavaScript ecosystem, a lack of dependencies may be a conscious decision to secure the supply chain. Motivations like these are the kind we aim to reveal with further work.
3.2. Case 2: The ‘classical’ package (11yr last updated)
Some packages don’t necessitate updates and unlike the previous example, may appear inactive; we call these ‘classical’ packages. The package imurmurhash [^4] https://github.com/jensyt/imurmurhash-js is an implementation of the MurmurHash3 hashing algorithm, last updated 11 years ago, with a single maintainer. Despite these factors, it still has 23.6 million users. There is only a single, closed issue concerning a false positive virus detection (dalerka, 2020); however, the maintainer still appears responsive.
One user in the issue suggests that based on the age of the package, and that the virus detection issue continues to exist, people should no longer use it in favour of more active packages.
Well, guys… Nothing has happened here in three years and the repo here hasn’t been updated in 10 (!) years. I would say it’s time to stop using this package and let everyone who uses it know about it.
The maintainer’s response, as of 2023, is simple:
The repo hasn’t been updated because there’s nothing to update, unfortunately.
This case exemplifies the ‘classical’ package case, as the implementation of an algorithm to a specification is unlikely to require the kinds of updates other packages do, especially when there are no dependencies to update. We use this example to illustrate that by reducing dependencies, developers may be able to reduce required maintenance.
3.3. Case 3: The trivial package (is-number)
Trivial packages are small NPM packages that implement a single task. Recent research has investigated the consequences of relying on these packages along the dependency chain (Abdalkareem et al., 2017). We look at the case of is-number [^5] https://github.com/jonschlinkert/is-number, one such trivial package with a simple task: checking if a value is a number. The package has 25.7 million users on GitHub, however it only has four contributors and hasn’t been updated in 6 years. This tiny package is part of the dependency chain of the popular eslint package. We present this case in contrast to the classical package case above.
With a number of open pull requests, an issue was opened urging the maintainer to merge some changes relating to optimisation, especially due to the popularity of the package (theUpsider, 2022).
This package has 69 million weekly downloads. Please merge the optimisation pull requests.
When one commenter asks why anyone even uses the package for such a simple task, another comments:
[The package] was in a package-json of mine. eslint-config-next uses is-number in a deeply nested dependency. ESLint might run 0.1% faster if the optimisation was implemented.
Whether the optimisation changes would be impactful or not, this package is clearly a case of the dependency chain relying on a trivial, inactive package with only a few contributors. Unlike the previous case of a classical package, however, there are possibly needed updates. The observation from is-number is that not every package with no dependencies is useful; in fact, is-number may be an obvious candidate for removal in projects looking to reduce dependencies themselves.
3.4. Case 4: Complex trivial package (internally vulnerable)
Some packages may appear trivial in that they solve a single problem, have a small project size or simple, one function API, typically meeting the requirements for a ‘trivial’ or ‘micro’ package, however, the task they deal with is deceptively complex. One common target area for software attacks is user input, and thus, functionality that manipulates strings or relies on regular expressions are often vectors to introduce vulnerabilities. packages that deal in these areas may be intrinsically more vulnerable to outside attack, necessitating more frequent updates compared to other cases of simple packages. The ms [^6] https://github.com/vercel/ms package is a package to convert string times to milliseconds. Like other trivial packages, this package is small (6.72kB) and performs a single task. However, a deeper look into the package reveals a different story: there are 36 contributors, 17 open issues and 12 pull requests, with the last commit 7 months ago. Additionally, Snyk reports two security vulnerabilities in past versions of the package: regular expression denial of service across two ranges of versions (Snyk, 2024a). Despite this, the package has 31.1 million users on GitHub.
This package is an example of how small, seemingly simple packages can introduce weak points in the dependency chain. In this case, the problems with this package stem from the complexities of regular expressions and string parsing. Due to the single functionality, lack of dependencies and small size, this package may appear deceptively trivial to potential users. The use of these types of packages may not necessarily be unwarranted though; by reusing a package, vulnerabilities can be reported, consumers of a package can be notified, and updates can be pushed.
3.5. Case 5: The bundled package (clone-and-own)
We present another interesting case, where a zero dependency package absorbed its dependencies. The chalk package [^7] https://github.com/chalk/chalk, for styling terminal text using ANSI, has no dependencies, but this was not always the case. Prior to removing all dependencies, the package had two dependencies: supports-color and ansi-styles, both also maintained under the GitHub chalk profile. In 2021, the packages were bundled; their files were moved into the vendor folder of the chalk package (sindresorhus, 2021), while the dependencies are still maintained separately. In the GitHub issue discussing the roadmap for this update, the main contributor explains their reasoning (sindresorhus, 2018):
The benefit is require speed and less dependencies, so faster installs. Usually, this would not be a problem, but Chalk is just used so much. Chalk is for Node.js. The consumer would never bundle themselves.
Multiple commenters expressed their concerns about bundling dependencies and its impact on the NPM ecosystem, for example:
[Bundling] has a number of downsides: No vulnerability reports upon install, no deprecation reports upon install (and no letting npm automatically select versions for you in the event something gets deprecated), guaranteed n * m install size were n is the number of different versions of Chalk within a large project (could be many, many times) and m is the number of dependencies Chalk has … I’m sure there are other downsides that I’m missing but it all feels like a hack and a move backwards. I’m not convinced either.
One user even calls into question if the dependencies should be separate to begin with:
If they are really not used by anyone else, why are they separate packages in the first place? I would find it concerning if packages start adopting this approach of bundling deps. It’s npm’s job to resolve dependencies, and only the top-level consumer should bundle. Some other concerns have been mentioned like what if there is a security vulnerability reported, and npm audit won’t report it, and nobody has the ability to upgrade through a soft range. What if the top-level consumer bundles, and only uses a fraction of chalk API (and transitive dependencies), but tree shaking cannot kick in because chalk and its deps are already bundled.
While the developer decided not to go through with bundling at the time, they later did so anyway, though further discussion cannot be found on the repository. Cases like these may help us understand how and when developers choose to reduce dependencies.
4. Research Agenda
Our early results reveal five cases of packages that do not follow the majority mindset of ‘reuse as much as possible’. As a side effect, these packages may in-turn be more resilient to package vulnerabilities and seem outdated, but are still heavily depended upon. We split our research agenda into five dimensions to further investigate how to utilize how these packages reuse dependencies.
4.1. Characterizing End-of-Chain packages: A New Lens for Reuse
The characterization of end-of-chain packages is crucial to understanding their role in the software ecosystem and the strategic decisions around their reuse. These packages, with minimal dependencies, provide unique insights into how reuse can be optimized for stability and resilience in software projects.
  1. (1)
    How can we characterize end-of-chain packages to better inform reuse decisions? Approach: Explore the unique characteristics of end-of-chain packages, such as their autonomy, low dependency footprint, and stability. Track the history to explore whether these packages once had dependencies that they eventually absorbed or removed.
  2. (2)
    What impact does removing or replacing end-of-chain packages have on the broader software ecosystem, and how should reuse strategies account for this? Approach: Investigate how the removal or replacement of end-of-chain packages affects projects that depend on them. We can study how upstream packages can be affected and whether the functionality can be easily absorbed by these upstream clients.
4.2. Strategic Dependency Management: When to Reuse or Not
To rethink reuse, it is essential to examine when and why dependencies should be reduced, absorbed, or replaced. Understanding the right timing and context for reusing packages or creating new solutions will optimize software performance, minimize risks, and improve long-term maintainability.
  1. (3)
    How can we optimize the decision-making process for when to reuse a package versus creating a new function? Approach: Research the trade-offs between using established packages and creating custom functions within the package.
  2. (4)
    What criteria should guide the decision to absorb (or bundle) dependencies into a package, and how can this inform the reuse strategy? Approach: Study the motivations behind bundling dependencies into a package (as seen with chalk) and how this affects the software ecosystem. Develop decision criteria that developers can use to decide whether to absorb or remove dependencies, taking into account long-term reuse and minimal dependency management.
4.3. Reimaging Supply Chain Resilience and Reuse Sustainability
The long-term sustainability and resilience of the software ecosystem depend on how packages are reused and maintained. Redefining reuse strategies that prioritize resilience and minimize risk can enhance the overall health of the ecosystem, especially in critical supply chains.
  1. (5)
    How can we redefine reuse as a strategy for building resilience and sustainability in the software supply chain? Approach: Redefine reuse not just as a cost-saving strategy, but as a means of enhancing the resilience and sustainability of software ecosystems. Examine how minimal dependency and end-of-chain packages can play a more central role in resilient supply chains, and propose guidelines for their reuse in critical systems.
  2. (6)
    What factors contribute to the long-term sustainability of packages at the end of the chain, and how can this inform reuse policies? Approach: Study the factors that enable the long-term sustainability of end-of-chain packages, such as their lack of dependencies, community support, and focus on stability. Develop best practices for maintaining and reusing such packages over extended periods, considering both their technical and community aspects.
  3. (7)
    What are the trade-offs between reuse and ecosystem health, and how can we ensure that the reuse of packages does not negatively impact the ecosystem? Approach: Examine the ecological effects of widespread package reuse, including the risks of over-reliance on certain packages and the potential for cascading failures. Propose a balanced approach to reuse that ensures packages are used in ways that contribute to ecosystem health without causing dependency bloat or security vulnerabilities.
4.4. Redefining the Future of Reuse in a Minimal Dependency Ecosystem
As end-of-chain packages become more prominent, it is crucial to rethink how reuse is performed in an ecosystem with minimal dependencies. This redefinition can lead to more efficient, secure, and sustainable software architectures in the long run.
  1. (8)
    Can we develop a new model for package reuse that prioritizes self-sufficiency and reduces upstream dependency risk? Approach: Propose a new model for package reuse that encourages self-sufficiency, reduces risk exposure from upstream vulnerabilities, and focuses on reducing unnecessary dependencies. Explore how end-of-chain packages can serve as examples for creating more autonomous software components in the ecosystem.
  2. (9)
    What role do end-of-chain packages play in shaping the reuse culture within open-source communities, and how can this inform future practices? Approach: Explore how end-of-chain packages influence reuse culture within open-source communities. Investigate whether they promote a more sustainable, self-sufficient model of software development, and provide recommendations for fostering a reuse culture that values long-term stability and security over short-term convenience.
4.5. On Rethinking Reuse for the Ecosystem
Further investigation into the nuances of reuse strategies is essential to fully understand how minimal dependency ecosystems can thrive. The following questions focus on measuring impact, governance, and creating feedback mechanisms that foster sustainable reuse.
  1. (10)
    How can we measure the impact of reduced dependency on both the quality and performance of reused packages? Approach: Develop performance and quality metrics for evaluating packages that reduce their dependencies. Study how reducing dependencies can improve performance, security, and maintainability, and create guidelines for balancing these metrics when deciding to reuse a package.
  2. (11)
    What are the governance implications of shifting to a minimal dependency, end-of-chain package model in terms of community support and project lifecycles? Approach: Investigate the governance implications of prioritizing minimal dependency models. How does this affect project lifecycles, developer participation, and long-term community support for these packages? Propose governance structures that encourage the adoption of minimal dependency practices.
  3. (12)
    How can we create a feedback loop where the successful use of end-of-chain packages leads to broader adoption of minimal dependency strategies? Approach: Examine case studies of projects that successfully adopted end-of-chain packages and explore how their success can be used as a model for broader adoption of minimal dependency strategies. Develop a framework for creating feedback loops that encourage the wider adoption of these practices.
5. Discussion and Conclusion
Our results suggest that end-of-chain packages potentially play an important role in the NPM ecosystem, with four of the ten most depended upon packages on the dependency chain having no dependencies themselves. While there are obvious benefits to a lack of dependence on other software, such as a lack of inherited vulnerabilities, other benefits remain unknown. Additionally, we don’t know how these gains stack up against the productivity benefits of code reuse. By looking at five example cases of packages with no dependencies at the end of the supply chain, we make the case that there may be new insights to be gained from similar packages across software ecosystems. End-of-chain packages may have their own risk and rewards that need to considered, and the motivations behind the use of no dependencies are currently unstudied. We present a research agenda for understanding these aspects of end-of-chain packages.
Answering the proposed research questions will take multiple different approaches. Firstly, the systematic analysis of a large dataset of packages across multiple ecosystems, including NPM, but also PyPI and Maven.As our current observations may not be generalisable to other software ecosystems, we hope to look beyond just NPM packages: different ecosystems have different policies and communities that could reveal different insights. To understand the motivations behind end-of-chain packages, we need to seek the perspectives of the developers behind these packages, via developer surveys, interviews or analysis of GitHub discussions. Ultimately, our hope is that by answering these questions, we can better understand when understanding of the concept of reuse.